evaluate:2cocec2300+3sin2600-3/4tan2300

Grade:

1 Answers

Yugabrat Gogoi
9 years ago
 please find your answer below-
 
2 cosec 230o + 3sin2600-3/4tan230o
=2(2)2+3[(31/2)/2]2 -3/4[1/(31/2)]
=8+(9/4) -(1/4)
=40/4
=10
 
hope you like this. thank you